OpenStudy (anonymous):

laplace transform y'+2y=e^-t, y(0)=1

OpenStudy (anonymous):

Take the Laplace transform of both sides: \[\mathscr{L}\left\{y'+2y\right\}=\mathscr{L}\left\{e^{-t}\right\}\] Recall \(\mathscr{L}\left\{y'\right\}=s\mathscr{L}\left\{y\right\}-y(0)\): \[s\mathscr{L}\left\{y\right\}-y(0)+2\mathscr{L}\left\{y\right\}=\mathscr{L}\left\{e^{-t}\right\}\\ (s+2)\mathscr{L}\left\{y\right\}-1=\frac{1}{s+1}\\ \mathscr{L}\left\{y\right\}=\frac{\frac{1}{s+1}+1}{s+2}\\ \mathscr{L}\left\{y\right\}=\frac{\frac{s+2}{s+1}}{s+2}\\ \mathscr{L}\left\{y\right\}=\frac{1}{s+1}\\ y=\mathscr{L}^{-1}\left\{\frac{1}{s+1}\right\}\]

OpenStudy (anonymous):

thank u